Supplemental Operations Report 20May2023

[title Supplemental Operations Report – May 20th]

Supplemental Operations Report 20May2023

Name of person filing report: Sergii Iakymov

Reason for Report: Routine.

Non-Nominal Systems: Robotic observatory.
Action taken for non-nominal systems: New parts for Robotic observatory are received and installation is pending.

Power system:
Solar: nominal. Input breaker of one of MPTT was off two times this week. After resetting the breaker and solar charger work was nominal until the next bright and hot day. This issue has to be investigated by a certified technician.SOC Last 24 hours: Max 100%; Min 46 %; Avg 71.2%.
VDC Last 24 hours: Max 58.04V; Min 44.34V; Avg 50.17V.Generator run time: not used. Generator hours 3810.3.

Propane Reading Station Tank: 75%
Propane Reading Director Tank: 59%
Propane Reading Intern Tank: 71%
Propane Reading Generator Tank: 72%

Water (Static Tank) – 339 gallons
Water in GreenHab – 165 gallons
Water (Outpost tank) – 324 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: n/a

Sojourner rover used: Yes.
Hours: 191.7
Beginning Charge: 100
Ending Charge: 100
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged.
ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: All parked at the outpost and awaiting to be taken to service.

HabCar used and why, where? Yes. To Hanksville for supplies to pick up.

Crew Car used and why, where? Yes. From Grand Junction by crew.

General notes and comments: Nominal.

Dual Split in Science Dome: off due to warmer weather.

Summary of Internet: Nominal

Summary of suits and radios: Nominal

EVA COMMS: Nominal

Campus wide inspection, if action taken, what and why: nothing to report

Summary of General Operations: All nominal.

Summary of Hab operations: Kitchen outlets are not working, inspection will be done after current crew rotation is over.

Summary of GreenHab Operations: Nominal, heater turned off.

Summary of SciDome Operations: Nominal.

Summary of any Observatory Issues: New parts for Robotic observatory are received and installation is pending.

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations:

Director’s trailer: previously reported leak in the bathroom sink pipe was actually located at the pipe-faucet connection; the connection was adjusted to resolve the issue.
The shed was coated a second time to finish outside panting.
The barrel for the burnable trash has been cleaned from ashes.

Summary of Health and Safety Issues: Nominal.

Questions, Concerns, Supplies needed and Requests: Nothing to report

Sol Summary – May 19th

Crew 281 Sol Summary Report 19 May 2023
Sol: 5
Summary Title: Rock sampling, radios and chocolate
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
EVA: Today’s EVA was a local MDRS campus surveillance. Crew of 2 did a visual inspection of the hab to ensure there were no safety concerns after yesterday’s storm, and there were none. They then walked over to go to the chosen site for the Pegasus scoop test. While assembling the scoop for the test, the battery connection was damaged and the test was aborted, this was later repaired in the hab by Ana (Crew Scientist) and KC (Crew Engineer). After video documenting the attempted Pegasus scoop test, Ritu (XO) did a drone survey of MDRS facility from 90 m altitude for the purpose of georeferencing the region for future cartographic aid.
In the Greenhab,and harvested microgreens, arugula and lettuce for a fresh salad lunch.
In the science dome, Megan (Commander) transplanted cacao seeds, Ana (Crew Scientist) worked on cartographic maps and other items.
In the RAM, retrieved supplies for projects.
For lunch Ritu (XO) made lentil stew with quinoa and Megan harvested the greens for salads. After the research and hab work, we did report writing. For dinner we ate lentils and rice along with mango chutney and tomato salsa.
Look Ahead Plan: Tomorrow we will continue scanline measurements at control point 3, Kissing Camel Ridge with a crew of 3. After that, another EVA for the antenna set up for the ham radio experiment will follow in the afternoon.
Anomalies in work: None
Weather: Nominal
Crew Physical Status: Nominal
EVA: Completed 1 today

Reports to be filed:
Greenhab Report
Ops Report
Journalist Report
Photo of the Day

Support Requested:
Mission support is requested in case of storms, if the antenna needs to come down or be retrieved.

Operations Report – May 19th

Crew 281 Operations Report 19-05-2023
SOL: 5
Name of person filing report: Megan Kane (Commander) and Ritupriya Patil (Executive Officer)
Non-nominal systems:
Electrical outlets in kitchen
Remote observatory
EVA Helmet
Notes on non-nominal systems: None
Spirit rover used: No
Hours: (before EVA): 222.2
Hours: (after EVA): NA
Beginning charge: (Before EVA): NA
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
Opportunity rover used: No
Hours: (before EVA): 129.6
Hours: (after EVA): NA
Beginning charge: (Before EVA): NA
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
Curiosity rover used: No
Hours: (before EVA): 239.8
Hours: (after EVA): NA
Beginning charge: (Before EVA): NA
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
Perseverance rover used: No
Hours: (before EVA): 264.3
Hours: (after EVA): NA
Beginning charge: (Before EVA): NA
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
General notes on rovers: no rovers were used
Summary of Hab operations: put notes here
WATER USE: 22.8 gallons
Water (static tank): 373.6 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet:
Summary of suits and radios: All suits and radios are fully recharged.
Summary of GreenHab operations:
WATER USE: 3.83 gallons
Heater: Off
Supplemental light: Off
Harvest: in grams :
Lettuce 34
Rocket (Arugula) 1
Microgreens 8
Summary of Science Dome operations:
Dual split: Heat and AC, Off
Cacao Experiment:
Transplanted cacao seeds
Geotechnical Experiment:
Worked on cartographic map and other items
Summary of RAM operations: (Tools used, work done)
Removed soft hammer for use during EVA to safely drive ground pin during antenna installation.
Summary of any observatory issues: NA
Summary of health and safety issues:
No new issues to report.
Questions, concerns and requests to Mission Support:
Request Mission Support assistance to take down or retrieve the ham radio antenna in case of emergency or high winds.

Journalist Report – May 19th

Crew 281 Journalist Report 19 May 2023

Journalist Report

By: Rachel Jones

Sol 5 and Still Alive!

The morning was mayhem and included a last-minute change in EVA plans to move my antenna set-up to Sol 6. Instead, Ritu and Ana had their hands full doing some exciting field tests using the robotic scoop and more drone aerial surveillance. Ana’s robotic scoop has been named “Pegasus” in honor of this mission (our crew is “Team Pegasus”), and there were some challenges with its initial deployment. During the deployment of the scoop, a pin broke off while connecting the battery. The lesson learned is that a human interface redesign is needed to allow easier field deployment by astronauts or those wearing gloves. Meanwhile, Ritu was able to capture some fantastic footage of the deployed scoop and also captured aerial images of the MDRS Campus while geo-referencing a control point for later EVAs. The last part of
the EVA involved an exterior inspection of the Hab and facilities. The EVA team investigated some of the sources of irregular banging heard during periods of high winds in the afternoon. No immediate threats were found, and the images provided should allow mission support to resolve the banging issue in the future.

Lunch today was stewed lentils and quinoa prepared by Ritu. It was served alongside a fresh salad of harvested lettuce, arugula, and microgreens from the GreenHab. Following lunch, the crew came together to perform afternoon chores and do routine maintenance. I imagine this part of the SIM greatly reflects reality, as there are always things to do. Unfortunately, most do not want to hear the trials and tribulations of checking the toilets and water systems.

After getting the day’s requirements taken care of, most of us went on to our individual projects. Ana helped Megan transplant the chocolate seeds into soil in the Science Dome. Ana, of course, started writing a few papers. Megan focused on future reports for MDRS. KC and I did a trial setup for the MDRS High Frequency (HF) antenna that we hope to deploy on tomorrow’s EVA. MDRS has a remarkably versatile amateur radio setup that was designed by Hope L. (ND2L) from the Spaceward Bound Utah’s first all high school which was out at MDRS before the 2022-2023 season began! KC also checked in on all the GreenHab plants and kept an eye on the power status in the mid-day heat. Ritu downloaded her drone footage and started on daily reports.

For dinner tonight, we have leftover lentils and fresh rice. I received some advice on troubleshooting the sound issues I’ve been having with my Software Defined Radio (SDR) set up and will spend the evening attempting various configurations. I’m looking forward to deploying my equipment tomorrow and hopefully getting on the air. For now, signing off from Mars.

GreenHab Report – May 19th

Crew 281 GreenHab Report 19 May 2023
GreenHab Officer: Megan Kane
Environmental control: mixed door open/ door open with fan
Average temperatures: 89.7 °F
Hours of supplemental light: None
Daily water usage for crops: 2.25 gallons
Daily water usage for research and/or other purposes: 1.59 gallons
Water in Blue Tank 168.51 gallons
Time(s) of watering for crops:
0700: refilled 7 water bottle (3.5 L) watered other plants as needed 2 L
1830: refilled 1 water bottle (500 mL) watered other plants as needed 2 L
Changes to crops:
Several of the tomatoes that had shown stress on Sol 0 have shed a few of the leaves that showed stress at the time. Their appearance had remained similar (chlorosis with purple veins and necrotic spots) for the past several days prior to abscision. New growth on all plants looks to be green and in good health.
Cucumbers are closer to being ready to harvest. There are 3 cucumbers and more flowers that may become cucumbers.
In the morning the GreenHab was checked and watered as required. Initial temperature was 71.4 degrees Fahrenheit at at 41 % humidity. Humidity remained above 10 % throughout the day. The cooler temperature and increased humidity seems to have cheered up the plants. While the greenhab made it up to 105 degrees, it did not stay there long due to the afternoon storm.
The watering experiment is going well.
The experimental cooling setup which uses a cloth with water seems to be helping a little with the temperature and humidity, espcially at night when the building can be closed, however we will know more in the next couple of days. The surface area for evaporation and amount of available water is limited but, with more matierials, we could potentially try and keep the temperature under 100 degrees Fahrenheit and increase humidity so that stomata can open to allow photosynthesis. The prototype passively evaporated 1 liter of water and was not refilled until evening. Tomorrow checks will be done earlier in the day to verify water level over time.
Harvest: (include which crop and mass in grams):
Lettuce 34
Rocket(Arugula) 1
Micro greens 8
Support/supplies needed: None

EVA Report – May 19th

Crew 281 EVA Report 19-05-2023

EVA # 6

Authors: Megan Kane (Commander), Ritupriya Patil (Executive Officer), Ana Pires (Crew Scientist)

Purpose of EVA:
Erect antenna to the North West of the hab’s rear airlock for Ham Radio set up (not achieved)
Test rock sampling device (Pegasus scoop) and collect samples.
External inspection of the hab, since there were some loose cables/bungee cords banging the hab walls during yesterday’s storms.
Video record the operations

Start time: 0838 hrs

End time: 0954 hrs

Shortly before the EVA it was determined we needed additional time to prepare and test the Ham Radio setup. It was decided that Rachel would focus on this and not go on EVA therefore the first purpose listed was not achieved.

At 0838 the two members of the EVA left. The first objective was to inspect all rovers and collect their information. Hours and State of Charge were recorded for each vehicle and it was verified that all were securely plugged in.

After this the EVA conducted a visual inspection of the hab to ensure there were no safety concerns after yesterday’s storm. Once that objective was complete, the EVA members went to the chosen test site for the Pegasus scoop test. While assembling the scoop for the test, the battery connection was damaged and the test was aborted. Ana (Crew Scientist) repaired it in the hab, post EVA. After documenting the attempted test and conducting a quick drone survey of the area the crew packed up the drone into its case. The drone was carried back to the hab and stored in Spirit’s cargo bay to be retrieved prior to return.

After a short water break they went to collect samples from the area surrounding the hab as described in purpose 2 of the EVA. They also took photographs of the area.

At 0928 the drone was deployed to take visual imagery of the habitat and surrounding area for geolocation mapping. Drone deployed to 80 m initially ascending to 97 m at 0933. The drone landed at 0939 with 42% battery remaining.

After a second watercheck the EVA crew made their way back to Spirit to collect the Pegasus Scoop. Enroute they discovered and retrieved some man made debris and retrieved it for disposal in the hab. They returned to the Hab entering the airlock at 0949.

Overview of EVA

Conduct Hab inspection
Test Pegasus Scoop
Collect samples
Document activities in photographs and video

Issues Experienced During EVA

While assembling the Pegasus Scoop for testing, the pin for the battery broke. The system was repair later in the day.
Discovered debris near hab


Crew completed Rover checks
Crew completed visual inspection of the habitat, with some pictures and videos
Crew Scientist conducted a test for deploying the Pegasus Scoop, a rock sampling device.
Conducted 2 drone flights to collect video footage of the Hab.
Collected 7 rock samples weighing 1068 g

Destination: MDRS Hab and surrounding 500 m radius
Coordinates (use UTM WSG 84): 0518100 E, 4250900 N
EVA Participants 2: Ritupriya Patil (Executive Officer), Ana Pires (Crew Scientist)
Road(s) and routes per MDRS Map: Walk around the Hab
Mode of travel: Walk

Copyright © The Mars Society. All rights reserved. | Main Site